# Word-specific tonal realizations in Mandarin

### v1

@@ 2.6.2 Cross validation
We fitted nine models to the training data. In addition to the eight models already introduced above, and assessed in Figure 4 (left-hand panel), we added one more model that was given data in which the values of word were randomly permuted. That is, tokens of a given word were now assigned different random word labels. In what follows, we refer to this model as the random-word model. If the effect of word is genuine, then random permutation of the word labels should significantly reduce prediction accuracy. To quantify model accuracy, we obtained the models’ predictions for the F0 contours of the held-out test data, and calculated the sum of squared errors (SSE) as measure of prediction accuracy. [^14] The sum of squared errors (SSE) is the sum of the squared difference between the observed and predicted values. A smaller SSE indicates more precise model predictions. The SSE for the held-out data of a given model should be smaller than the SSE of the baseline model if the addition of one or more predictors indeed improves that model’s prediction accuracy. We implemented cross-validation using both the scaled t-distribution (scat) and simple Gaussian models. As models with scat are computationally expensive to estimate, for environmental reasons we used these models for one held-out dataset only and compared the results with those of less expensive Gaussian models for the same held-out data. Since prediction accuracy turned out to pattern similarly for scat and standard Gaussian models, and since fitting Gaussian models requires much less computational resource, we ran Gaussian models to cross-validate our results for 100 random held-out datasets.
@@
Cross-validation results are presented in Figure 6. The left-hand panel presents the SSE difference between the baseline model and each of the nine models of interest, for one random held-out dataset. Larger positive values indicate that the model of interest offers more precise predictions than the baseline. The difference in SSE obtained with a standard Gaussian model is shown in blue, and the difference obtained with a scat model is shown in orange. All the individual segment-related controls increase prediction accuracy over the baseline, albeit to varying degrees. However, the omnibus-segment model and the word model produce significantly greater increases in prediction accuracy, with the latter reducing the SSE to a larger extent than the former, replicating the model fit results. Moreover, when word labels are randomized, model accuracy plummets: the SSE of the random-word model is greater than that of the baseline model.
Because the results of GAM models with and without scat are very similar for the first set of test data, we infer that the distribution of results across the 100 cross-validation runs using Gaussian models will also be informative. These results are reported in the right-hand panel of Figure 6. The overall pattern is very similar to the result of the initial single cross-validation run. The word model is still far more accurate than all the other models, and significantly outperforms the omnibus-segment model. In other words, the greater reduction in AIC that we observed for the word model in Section 2.6.1 is not just a side-effect of an evaluation procedure that happens to favor that model over the omnibus-segment model. The word model truly provides enhanced predictions.
@@
The results presented so far provide strong evidence that word type is predictive of tonal realization, over and above the segment-related predictors established by previous studies. Our hypothesis, arising from the theoretical framework of the Discriminative Lexicon Model (DLM Baayen et al., , 2019; Chuang and Baayen, , 2021; Heitmeier et al., 2023c, ), is that the predictive power of word type arises not only from articulatory constraints, but also from a close association between word meaning and phonetic form, which enables the language learner or user to discriminate more efficiently between forms with different meanings. However, since word meaning varies with context, there is no one-to-one correspondence between word type and the meaning of a given token; word can only encompass a rather general approximation of what each token means. If word meaning is indeed predictive of tonal realization, then a model replacing word by sense should improve model fit even further. [^15] Note, however, that as discussed in Section 2.3, the senses that constitute the possible values of our sense variable discretize a much more subtle and interesting palette of shades of meanings.
Figure 6: Model accuracy under cross validation for the word analysis. The left-hand panel presents the reduction in the sum of squared errors (SSE) for models using one set of test data. Positive values indicate lower SSE as compared to the SSE of the baseline model. Orange dots represent scat models, while blue dots represent Gaussian models. The right-hand panel shows the results of 100 runs of cross validation, using Gaussian models. The boxplots represent the distributions of reduction in SSE.
@@ 2.7.1 Evaluation of predictors
Figure 7 shows model fit improvement and concurvity for models based on the smaller dataset that included at least 14 tokens of each word sense. Even with this smaller dataset, the overall pattern of results is very similar to that of the word type analysis. Critically, however, sense appears to be a somewhat better predictor than word. Not only does it account for more variance in F0, resulting in better model fit, but its effect is also less confounded with other covariates in the model, as indicated by a smaller concurvity score.
@@
Figure 7: Sense analysis. The left-hand panel shows model fit improvement, gauged by decrease in AIC units when a predictor (or set of predictors) is added to the baseline model. The right-hand panel shows the concurvity score of individual predictors in three models: the omnibus-segment model with all segment-related control variables (blue), the word-type model with word as predictor (red), and the sense model with sense as predictor (purple).
@@
Figure 8 presents the predicted tonal contours for different senses of three words: {CJK*}UTF8bsmi不要 búyào (left), {CJK*}UTF8bsmi實在 shízài (upper right), and {CJK*}UTF8bsmi能夠 nénggòu (lower right). The word {CJK*}UTF8bsmi不要 búyào is a polysemous negation marker in Mandarin. The four senses that are found in our dataset are ‘prohibition’, ‘dissuasion’, ‘unneccesity’, and ‘to wish something to not happen’ (s1 to s4, respectively). It can be seen that the different senses have clearly different tonal realizations. The panels on the right-hand side of Figure 8 present the predicted contours for the other two words, each of which has two senses in our data. For {CJK*}UTF8bsmi實在 shízài, tonal realizations vary greatly between the two senses (‘truly’ and ‘indeed’), whereas the realizations of the two senses of {CJK*}UTF8bsmi能夠 nénggòu (‘being capable’ and ‘enabling’) are more alike, and differ mainly with respect to the amplitude of the pitch inflection.
@@
Figure 8: Sense analysis. Examples of the pitch contours predicted by the general smooth for time for female speakers, combined with the partial effects of the factor smooth for sense. The left-hand panel shows the fitted tonal contours for different senses of the word {CJK*}UTF8bsmi不要 búyào, a negation marker in Mandarin. The four senses are ‘prohibition’, ‘dissuasion’, ‘unneccesity’, and ‘to wish something to not happen’. The upper right-hand panel shows the fitted tonal contours for the two senses of {CJK*}UTF8bsmi實在 shízài, meaning ‘truly’ and ‘indeed’ respectively. The lower right-hand panel plots the fitted contours for the two senses of {CJK*}UTF8bsmi能夠 nénggòu: ‘being capable of’ and ‘enabling’.
2.7.2 Cross validation
@@
As shown in the left-hand panel of Figure 9, for the first random held-out dataset the model that has access to sense is more accurate than the model that has access to word, irrespective of the type of model used (scat or Gaussian). However, in 100 Gaussian cross-validation runs, it turns out that the model with sense is not necessarily always more accurate than the model with word (right-hand panel of Figure 9). Although the median of the reduction in SSE for the sense model is slightly higher than that for the word-type model, the variance of the sense model is somewhat larger.
Figure 9: Model accuracy under cross validation for the sense analysis. The left-hand panel presents the reduction in summed squared errors (SSE) for models using one set of test data. Positive values indicate lower SSE as compared to the SSE of the baseline model. Orange dots represent scat models, while blue dots represent Gaussian models. The right-hand panel shows the results of 100 runs of cross validation, using Gaussian models. The boxplots represent the distributions of reduction in SSE.
@@
Figure 10: Sense analysis. Predicted pitch contours of the partial effects of the factor smooth for sense, for the five most frequent senses (upper row) and the least frequent senses (lower row). Numbers in the parentheses indicate the number of tokens in the dataset for the different senses. Confidence intervals are anti-conservative, due to autocorrelation in the errors.
2.8 Random forest analyses
The analyses presented thus far were all conducted with GAMs. To make sure that the effects of word and sense are not method-specific, and to further explore the relative importance of these variables in predicting F0 contours, we also made use of random forests, a machine learning algorithm that is based on multiple recursive partitioning decision trees (Breiman, , 2001). Since this method focuses on overall prediction accuracy, we can include all predictors in one model, without needing to worry about concurvity or model interpretability. Figure 11 shows the variable importance scores for all our predictors in two random forest analyses carried out with the cforest implementation in the party package for R (Hothorn et al., , 2006), using the same dataset as the sense-based GAM.
The left-hand panel of Figure 11 shows the results of a random forest analysis conducted straightforwardly on log F0 values. Perhaps unsurprisingly, speaker and gender emerge as the most important predictors; sense is ranked third and, in line with our hypothesis, contributes to model accuracy more than word does. However, it is surprising that among all the predictors in this model normalized time receives a rather low ranking, given that the fall-rise-fall pattern is visible across the time series of many of the tokens in the dataset. Combined with the high importance of speaker and gender, this suggests that the algorithm is picking up more on the overall pitch level of tokens than on pitch variation within tokens. Furthermore, it is conceivable that individual word meanings have their own emotional valence that gives rise to higher F0 for positive words and lower F0 for negative words (Belyk and Brown, , 2014).
To sidestep the confounds described in the previous paragraph, we centered the pitch contours for each speech token separately by subtracting the mean pitch of its time series from each of its pitch values. The results obtained when a random forest is requested to predict centered pitch are presented in the right-hand panel of Figure 11. Now, normalized time is the most important predictor, whereas gender has become much less important. However, speaker remains high in the ranking, suggesting that speaker-specificity in tonal realization is substantial. Crucially, sense again emerges as a more important predictor than word.
Figure 11: Variable importance scores for each predictor in the random forest analyses for log pitch (left) and centered log pitch (right).
@@ 2.9 Summary of Section 2
The results presented in this section have provided evidence in support of our first two predictions. Firstly, variation in tonal realization cannot be reduced to the segment-related constraints on articulation previously described in the literature. Secondly, information about a word’s meaning in context improves prediction of its tonal realization in that context. To the extent that sense labels provide more fine-grained meaning distinctions than word labels do, our results suggest that meaning plays a role in shaping the realization of tonal contours in Mandarin. In other words, in addition to the relevant segmental differences previously identified in the literature, differences in meaning also contribute.
Nevertheless, as discussed in Section 2.3, sense labels impose discrete categories on semantic variation that is actually much richer, more subtle and more nebulous than can be captured by such inventories. In the computational models reported in Section 3 below, this problem did not arise since our use of the DLM enabled us to replace relatively crude sense categories with token specific semantic representations.
@@ 3 Understanding and producing item-specific F0 contours
So far we have shown that it is possible to identify meaning-specific modulations of the pitch contour for Mandarin words with the RF tonal pattern. The question therefore arises as to whether native speakers of Mandarin can in principle profit from these meaning-specific modulations. In other words, are the semantic components in words’ pitch contours sufficiently informative that they could facilitate word comprehension for the listener? A related question is whether these subtle semantic modulations are learnable for the speaker, as opposed to arising mechanically each time a word is produced, as one might expect for purely articulatory effects. As outlined in Section 1, our third and fourth predictions, repeated here for convenience, anticipate an affirmative answer to both these questions:
@@
    Given a pitch contour, the meaning of its carrier token can be predicted above chance level, assuming the listener has previous experience of that word type.
@@
    Assuming they have previous experience of a given word type, a speaker can produce an appropriate pitch contour for a meaning they want to convey with that word.
In this section of the paper, we explore these predictions with computational modeling using the Discriminative Lexicon Model (DLM Baayen et al., , 2019; Chuang and Baayen, , 2021; Heitmeier et al., 2023c, ). If we can show that a simple computational model can learn to predict the meaning of a word token from its pitch contour, and that pitch contours can be predicted from intended meaning, we have a proof of concept for the functionality of meaning-specific pitch realization in human lexical processing.
As described in Section 1, the DLM focuses on the relationship between words’ forms and their meanings, and allows for fine-grained alignments between low-level features of form and low-level features of meaning. Form-meaning relationships are captured by two networks: a comprehension network that maps word form onto word meaning, and a production network that maps word meaning onto word form. Recall that in the DLM theory of the mental lexicon, forms and meanings do not have representations in memory. Form representations represent ephemeral auditory or visual input, which dynamically generates a corresponding, equally ephemeral, meaning representation. Conversely, a meaning conceptualized by a speaker at a given point in time is dynamically transformed into ephemeral representations driving articulation. In line with this theory, the DLM generates forms and meanings on the fly on a token by token basis, making it is possible to model the relationship between a given token’s specific pitch contour and that token’s context-specific meaning, and hence to account for correspondences between meaning and fine phonetic detail. [^16] This is not possible in models of speech production and comprehension that rely on stored abstract representations to mediate between form and meaning, where every token of a given word type is assumed to be associated with the same stored representations (e.g., Cutler and Clifton Jr., , 1999; Levelt et al., , 1999).
In the DLM, both the form and the meaning of each word token are operationalized mathematically as high-dimensional numeric vectors. In order to test our predictions about the functionality of tonal modulations, the form vectors used in this study are based exclusively on the F0 contour of the relevant token, with no information about its segmental make-up. The meaning vectors that we use are context-specific, and hence also vary from token to token. Sections 3.1 and 3.2 describe how we obtained the vectors for meaning and pitch, respectively; Section 3.3 addresses the functionality of pitch in comprehension and 3.4 does the same for production.
@@ 3.1 Representing meaning: contextualized embeddings
Embeddings are widely-used numeric representations of words’ meanings, developed from the distributional semantic insight that words with similar meanings tend to occur in similar contexts (Firth, , 1968; Harris, , 1954; Salton et al., , 1975; Landauer and Dumais, , 1997). Embeddings represent word meanings as real-valued high-dimensional vectors in a semantic space (Schütze, , 1992). First generation word embeddings are static, type-level representations that model the meaning of a word type as a fixed point in semantic space, regardless of its usage in a given context. These representations therefore have difficulty distinguishing between multiple senses of a word (Pilehvar and Camacho-Collados, , 2020), and although various methods have been proposed to incorporate sense or context information into type-level embeddings (see e.g., Reisinger and Mooney, , 2010; Huang et al., , 2012; Neelakantan et al., , 2014; Iacobacci et al., , 2015), most of these methods involve the use of sense-annotated corpora, which as far as we know are not available off the shelf for Mandarin and, in any case, have the disadvantage of discretizing more complex semantic variability. An alternative is to use contextualized embeddings (CEs). In contrast to static, type-level embeddings, which are based on word co-occurrences irrespective of order, CEs take into account the sequence of words in the immediate context of a target word. CEs therefore encode word meanings at the token level, and different tokens of the same word type will have different but similar context-specific embeddings (see Appendix B for an informal introduction).
@@
To address the issue of words having context-specific meanings, this study used CEs produced for the tokens of our data by a pre-trained unidirectional language model based on the GPT-2 architecture. The model, developed by CKIP, Academia Sinica, Taiwan [^17] We use ckiplab/gpt2-base-chinese, which is available on https://github.com/ckiplab/ckip-transformers., was trained on a 4.3 billion character dataset written with traditional Chinese characters. The model has 102 million parameters and encodes each character as a 768-dimensional vector. We presented the target words and their preceding contexts (consisting of all the words that occur before the target in the current utterance, as well as all the words in the immediately preceding utterance) to the GPT-2 model, and obtained two embeddings from the model, one for each character. We then averaged the two embeddings, so that every token in our dataset received a 768-dimensional vector representing its context-specific meaning.
@@
To visualize the semantic space of the CEs, we reduced the 768-dimensional semantic space to two dimensions using tSNE (Van der Maaten and Hinton, , 2008). Figure 12 shows the resulting reduced 2D plane, using convex hulls to highlight that tokens of different word types typically fall in distinct regions while tokens of the same word type form clear clusters. Although perhaps unsurprising, this distribution confirms that, despite polysemy, the word types in our data do capture a general approximation of what each token means. It is also reassuring to see that, like static embeddings, the CEs can capture inter-word semantic relations. For instance, there is a cluster of school-related words in the lower left: {CJK*}UTF8bsmi學校 xuéxiào ‘school’, {CJK*}UTF8bsmi研究 yánjiù ‘research’, {CJK*}UTF8bsmi學到 xuédào ‘learn+resultative’ (see Vulić et al., , 2020, for similar results).

### v2

@@ 2.6.2 Cross-validation
We fitted ten models to the training data. In addition to the baseline model and the eight models assessed in Figure 4 (left-hand panel), we added one more model that was given data in which the values of word were randomly permuted. That is, tokens of a given word were now assigned different random word labels. In what follows, we refer to this model as the random-word model. If the effect of word is genuine, then random permutation of the word labels should substantially reduce prediction accuracy. To quantify model accuracy, we obtained the models’ predictions for the F0 contours of the held-out test data, and calculated the sum of squared errors (SSE) as a measure of prediction accuracy. [^22] The sum of squared errors (SSE) is the sum of the squared difference between the observed and predicted values. A smaller SSE indicates more precise model predictions. The SSE for the held-out data of a given model should be smaller than the SSE of the baseline model if the addition of one or more predictors indeed improves that model’s prediction accuracy. We ran all ten models with all 100 random splits to cross-validate our results for the 100 held-out datasets.
@@
Figure 6 presents boxplots of the SSE difference between the baseline model and each of the nine models of interest. Positive values indicate that the model in question offers more precise predictions than the baseline, as its SSE is smaller than the baseline’s. All the individual segment-related controls increase prediction accuracy over the baseline to some extent, albeit to varying degrees. However, the omnibus-segment model and the word model produce substantially greater increases in prediction accuracy, with the latter reducing the SSE to a larger extent than the former, replicating the model fit results. Moreover, when word labels are randomized, model accuracy plummets: the SSE of the random-word model is greater than that of the baseline model.
@@
The results presented so far provide strong evidence that word type is predictive of tonal realization, over and above the segment-related predictors established by previous studies. Our hypothesis, arising from the theoretical framework of the Discriminative Lexicon Model (DLM Baayen et al.,, 2019; Chuang and Baayen,, 2021; Heitmeier et al.,, 2025), is that the predictive power of word type arises not only from articulatory constraints, but also from a close association between word meaning and phonetic form, which enables the language learner or user to discriminate more efficiently between forms with different meanings. However, as discussed in Section 2.4, the effect of word type cannot be unequivocally attributed to semantics, since in addition to word meaning, word captures all of a word’s form properties, not only those previously identified as affecting tonal realization. Furthermore, since word meaning varies with context, there is no one-to-one correspondence between word type and the meaning of a given token; word can only encompass a rather general approximation of what each token means. To address both these issues, it is necessary to turn to Prediction 2 and the model with a factor smooth for sense as sole predictor added to the baseline. If word meaning is indeed predictive of tonal realization, then a model replacing the factor smooth for word by a factor smooth for sense should improve model fit even further. [^23] Note, however, that as discussed in Section 2.3, the senses that constitute the possible values of our sense variable discretize a much more subtle and interesting palette of shades of meanings.
Figure 6: Model accuracy under 100 runs of cross-validation for the word-type analysis. The boxplots represent the distributions of reduction in SSE.
@@ 2.7.1 Evaluation of predictors
Figure 7 shows model fit improvement and concurvity for models based on the smaller dataset that included at least 14 tokens of each word sense. Even with this smaller dataset, the overall pattern of results is very similar to that of the word type analysis. Critically, however, sense appears to be a somewhat better predictor than word. For this dataset, adding only a factor smooth for sense to the baseline model results in a fall of 6,443 AIC units compared with a fall of 6,078 AIC units when only a factor smooth for word is added. The difference of 365 AIC units means that the sense model is 1.81e+79 times more likely than the word model to explain the observed data. Furthermore, the effect of sense is also less confounded with the other predictors in the model (i.e., the baseline controls), as indicated by a smaller concurvity score.
@@
Figure 7: The left-hand panel shows model fit improvement gauged by decrease in AIC units when a predictor (or set of predictors) is added to the baseline model for the sense analysis. The right-hand panel shows the concurvity score of individual predictors in three models using the smaller dataset of 3,458 tokens: the omnibus-segment model with factor smooths for all segment-related control variables (blue), the word-type model with a factor smooth for word (red), and the sense model with a factor smooth for sense (purple).
@@
Figure 8 presents the predicted tonal contours for different senses of three words: {CJK*}UTF8bsmi不要 bu2yao4 (left), {CJK*}UTF8bsmi實在 shi2zai4 (upper right), and {CJK*}UTF8bsmi能夠 neng2gou4 (lower right). The word {CJK*}UTF8bsmi不要 bu2yao4 is a polysemous negation marker in Mandarin. The four senses that are found in our dataset are ‘prohibition’, ‘dissuasion’, ‘unneccesity’, and ‘to wish something to not happen’ (s1 to s4, respectively). It can be seen that the different senses have clearly different tonal realizations. The panels on the right-hand side of Figure 8 present the predicted contours for the other two words, each of which has two senses in our data. For {CJK*}UTF8bsmi實在 shi2zai4, tonal realizations vary greatly between the two senses (‘truly’ and ‘indeed’), whereas the realizations of the two senses of {CJK*}UTF8bsmi能夠 neng2gou4 (‘being capable of’ and ‘enabling’) are more alike, and differ mainly with respect to the amplitude of the pitch inflection.
@@
Figure 8: Examples of the pitch contours predicted by the general smooth for time for female speakers, combined with the partial effects of the factor smooth for sense. The left-hand panel shows the fitted tonal contours for different senses of the word {CJK*}UTF8bsmi不要 bu2yao4, a negation marker in Mandarin. The four senses are ‘prohibition’, ‘dissuasion’, ‘unneccesity’, and ‘to wish something to not happen’. The upper right-hand panel shows the fitted tonal contours for the two senses of {CJK*}UTF8bsmi實在 shi2zai4, meaning ‘truly’ and ‘indeed’ respectively. The lower right-hand panel plots the fitted contours for the two senses of {CJK*}UTF8bsmi能夠 neng2gou4: ‘being capable of’ and ‘enabling’.
2.7.2 Cross-validation
@@
As shown in Figure 9, for the 100 cross-validation runs, it turns out that the model with sense is not necessarily always more accurate than the model with word. The medians of the reduction in SSE for the word and the sense model are very similar, with the variance of the sense model being somewhat larger.
Figure 9: Model accuracy under 100 runs of cross-validation for the sense analysis. The boxplots represent the distributions of reduction in SSE.
@@
Figure 10: Predicted pitch contours of the partial effects of the factor smooth for sense, for the five most frequent senses (upper row) and the five least frequent senses (lower row). Numbers in parentheses indicate the number of tokens in the dataset for the different senses.
@@ 2.8 Summary of Section 2
The results presented in this section have provided evidence in support of our first two predictions. Firstly, word type is a stronger predictor of tonal realization than all the previously established word-form related predictors combined. Secondly, information about a word’s meaning in context improves prediction of its tonal realization in that context. To the extent that sense labels provide more fine-grained meaning distinctions than word labels do, our results suggest that meaning plays a role in shaping the realization of tonal contours in Mandarin. In other words, in addition to the relevant segmental differences previously identified in the literature, differences in meaning also contribute.
Nevertheless, as discussed in Section 2.3, sense labels impose discrete categories on semantic variation that is actually much richer, more subtle and more nebulous than can be captured by such inventories. In the computational models reported in Section 3 below, this problem did not arise since our use of the DLM enabled us to replace relatively crude sense categories with token-specific semantic representations.
@@ 3 Understanding and producing item-specific F0 contours
So far we have shown that it is possible to identify meaning-specific modulations of the pitch contour for Mandarin words with the RF tonal pattern. The question therefore arises as to whether native speakers of Mandarin could in principle profit from these meaning-specific modulations. In other words, are the semantic components in words’ pitch contours sufficiently informative that they could facilitate word comprehension for the listener? A related question is whether these subtle semantic modulations are learnable for a speaker, as opposed to arising mechanically each time a word is produced, as one might expect for purely articulatory effects. As outlined in Section 1, our third and fourth predictions, repeated here for convenience, anticipate an affirmative answer to both these questions:
@@
    Given a pitch contour, the meaning of its carrier token can be predicted above chance level by a simple computational model with previous experience of that word type.
@@
    Assuming it has previous experience of the relevant word type, a simple computational model can produce an appropriate pitch contour for a given meaning.
In this section of the paper, we explore these predictions with computational modeling using the Discriminative Lexicon Model (DLM, Baayen et al.,, 2019; Chuang and Baayen,, 2021; Heitmeier et al.,, 2025). If we can show that a simple computational model can learn to predict the meaning of a word token from its pitch contour, and that pitch contours can be predicted from intended meaning, we have a proof of concept for the potential functionality of meaning-specific pitch realization in human lexical processing.
As described in Section 1, the DLM focuses on the relationship between words’ forms and their meanings, and allows for fine-grained alignments between low-level features of form and low-level features of meaning. Form-meaning relationships are captured by two networks: a comprehension network that maps word form onto word meaning, and a production network that maps word meaning onto word form. Recall that in the DLM theory of the mental lexicon, forms and meanings do not have representations in memory. Form representations represent ephemeral auditory or visual input, which dynamically generates a corresponding, equally ephemeral, meaning representation. Conversely, a meaning conceptualized by a speaker at a given point in time is dynamically transformed into ephemeral representations driving articulation. In line with this theory, the DLM generates forms and meanings on the fly on a token by token basis, making it possible to model the relationship between a given token’s specific pitch contour and that token’s context-specific meaning, and hence to account for correspondences between meaning and fine phonetic detail. [^24] This is not possible in models of speech production and comprehension that rely on stored abstract representations to mediate between form and meaning, where every token of a given word type is assumed to be associated with the same stored representations (e.g., Cutler and Clifton Jr.,, 1999; Levelt et al.,, 1999).
In the DLM, both the form and the meaning of each word token are operationalized mathematically as high-dimensional numeric vectors. In order to test our predictions about the potential functionality of tonal modulations, the form vectors used in this study are based on the F0 contour of the relevant token. The meaning vectors that we use are context-specific, and hence also vary from token to token. Sections 3.1 and 3.2 describe how we obtained the vectors for meaning and pitch, respectively; Section 3.3 addresses the functionality of pitch in comprehension and 3.4 does the same for production.
@@ 3.1 Representing meaning: contextualized embeddings
Embeddings are widely-used numeric representations of words’ meanings, developed from the distributional semantic insight that words with similar meanings tend to occur in similar contexts (Firth,, 1968; Harris,, 1954; Salton et al.,, 1975). Embeddings represent word meanings as real-valued high-dimensional vectors in a semantic space (Schütze,, 1992). They have been found to provide a plethora of novel insights in both psychology (Landauer and Dumais,, 1997; Günther et al.,, 2019; Bruni et al.,, 2014) and linguistics (Gahl and Baayen,, 2024; Perek and Hilpert,, 2017; Nieder et al.,, 2023), and are widely used in computer science (Bojanowski et al.,, 2017). First generation word embeddings are static, type-level representations that model the meaning of a word type as a fixed point in semantic space, regardless of its usage in a given context. These representations therefore have difficulty distinguishing between multiple senses of a word (Pilehvar and Camacho-Collados,, 2020), and although various methods have been proposed to incorporate sense or context information into type-level embeddings (see e.g., Reisinger and Mooney,, 2010; Huang et al.,, 2012; Neelakantan et al.,, 2014; Iacobacci et al.,, 2015), most of these methods involve the use of sense-annotated corpora, which as far as we know are not available off the shelf for Mandarin and, in any case, have the disadvantage of discretizing more complex semantic variability. An alternative is to use contextualized embeddings (CEs). In contrast to static, type-level embeddings, which are based on word co-occurrences irrespective of order, CEs take into account the sequence of words in the immediate context of a target word. CEs therefore encode word meanings at the token level, and different tokens of the same word type will have different but similar context-specific embeddings.
@@
To address the issue of words having context-specific meanings, this study used CEs produced for the tokens of our data by a pre-trained unidirectional language model based on the GPT-2 architecture. The model, developed by CKIP, Academia Sinica, Taiwan [^25] We used ckiplab/gpt2-base-chinese, which is available on https://github.com/ckiplab/ckip-transformers., was trained on a 4.3 billion character dataset written with traditional Chinese characters. The model has 102 million parameters and encodes each character as a 768-dimensional vector. We presented the target words and their preceding contexts (consisting of all the words that occur before the target in the current utterance, as well as all the words in the immediately preceding utterance) to the GPT-2 model, and obtained two embeddings from the model, one for each character. Following standard machine-learning practice (e.g., Huang et al.,, 2021) we then averaged the two embeddings, so that every token in our dataset received a 768-dimensional vector representing its context-specific meaning.
@@
To visualize the semantic space of the CEs, we reduced the 768-dimensional semantic space to two dimensions using tSNE (Maaten and Hinton,, 2008). Figure 11 shows the resulting reduced 2D plane, using convex hulls to highlight that tokens of different word types typically fall in distinct regions while tokens of the same word type form clear clusters. Although perhaps unsurprising, this distribution confirms that, despite polysemy, the word types in our data do capture a general approximation of what each token means. It is also reassuring to see that, like static embeddings, the CEs can capture inter-word semantic relations. For instance, there is a cluster of school-related words in the lower left: {CJK*}UTF8bsmi學校 xue2xiao4 ‘school’, {CJK*}UTF8bsmi研究 yan2jiu4 ‘research’, {CJK*}UTF8bsmi學到 xue2dao4 ‘learn+resultative’ (see Vulić et al.,, 2020, for similar results).